OpenStudy (anonymous):

find the values of a,b and c that make the function \[f(r) = ar^{2} + br + c; \] \[r<0 , 2\sin^{2}(r) + 3\cos^2(r)\] \[r \ge 0\] twice differentiable at r=0.

OpenStudy (anonymous):

ok ready?

OpenStudy (anonymous):

^.^i already pasted it to a notebook xD let me delete the old post

OpenStudy (anonymous):

ok yes i'm ready lets go

OpenStudy (anonymous):

no damn!

OpenStudy (anonymous):

o.O ?

OpenStudy (anonymous):

let me say it in english

OpenStudy (anonymous):

for this to be continuous, the limit from the left must be equal to the limit from the right at 0

OpenStudy (anonymous):

the limit from the right at 0 is 3 the limit from the left at 0 is C this means C = 3

OpenStudy (anonymous):

hmm ahh yea i see!

OpenStudy (anonymous):

now i will try to write it in math \[\lim_{r\rightarrow 0^-}f(r)=\lim_{r\rightarrow 0^-} ar^2+br+c=c\] \[\lim_{r\rightarrow 0^+}f(r)=\lim_{r\rightarrow 0^+}2\sin^2(r)+3\cos^2(r)=2\sin^2(0)+3\cos^3(0)=3\]

OpenStudy (anonymous):

so we know C = 3

OpenStudy (anonymous):

AHH WoW, now i see man

OpenStudy (anonymous):

now to b \[f'_-(r)=2a+b\] \[f'_+(r)=2\cos(r)\sin(r)-6\cos(r)\sin(r)\] so \[f'_-(0)=b\] and \[f'_+(0)=0\] so \[b=0\]

OpenStudy (anonymous):

and finally \[f''_-(r)=2a\] \[f''_+(r)=-2\cos(2r)\] so \[f''_-(0)=2a\] and \[f''_+(0)=-2\] thus \[2a=-2\iff a=-1\]

OpenStudy (anonymous):

ahh yea

OpenStudy (anonymous):

really good job :D

OpenStudy (anonymous):

and your function is \[f_-(r)=-r^2+3\]

OpenStudy (anonymous):

yeah well it comes with no guarantee, but i believe it is correct

OpenStudy (anonymous):

sorry it took so long to write fingers are slow today

OpenStudy (anonymous):

np man :D

OpenStudy (anonymous):

and thanks much

OpenStudy (anonymous):

btw this line \[f'_+(r)=2\cos(r)\sin(r)-6\cos(r)\sin(r)\] was a mistake, it should have been \[f'_+(r)=4\cos(r)\sin(r)-6\cos(r)\sin(r)\] but you still get 0 at 0 so no harm done

OpenStudy (anonymous):

and the second derivative is correct. i just typed it wrong here

OpenStudy (anonymous):

:D oka :) Really gj

OpenStudy (anonymous):

cool ......lol

OpenStudy (anonymous):

:D can you have a look at the other question which is new :)